INAF Pisces observing, 2011 November 28-29

Observer: C. Arcidiacono, G. Brusa, K. Boutsia, M. Fumana
Telescope Operator: D. Gonzales-Huerta
Support Astronomer: O. Kuhn
Instrument: Pisces

Summary

Thin clouds on the sky in the afternoon. We kept having clouds for the whole night and it got worse during the second half. We had 1.5-3mag of extinction.
We open up early (4:00PM) to take narrow band flats with Pisces. Many flat frames are saturated due to cloud reflections. We obtained high and low count flats for Br-g, FeII, K, and some J.
The seeing has been <1.0'' for most of the night but it was very variable with peaks of 2.0'' giving trouble to the AO system. There is a plot of the Seeing vs. UTtime in /home/INAF/ScienceRuns/Nov2011/.
We lost 1h30min at the beginning of the night due to problems with starting up the AO system. After calling Alfio (in Italy), the problem was identified and it was solved by restarting the AO arbitrator.
For the rest of night we had many failed presets and shell rips, but most of them were caused because of variable seeing and high extinction because of clouds. The AO could not keep up with such variable conditions.
Neveretheless we have been able to take data for all targets scheduled for this night and the T-Tau project (that was the brightest target) is considered complete.

For more technical details see log in the AO twiki by G. Brusa at Log_20111129 .

Data Summary

Project Objects Filter ExpTime Notes
Bono IC10 Ks 20min (+10min sky) it was cloudy and AO kept failing - maybe some useful data
Palla T-Tau 2.14, Br-g, H2, FeII 1hr30min it should be 30minXfilter - passing clouds for part of the dataset - 3mag of extinction
Greggio NGC1560 Ks , J 30min, 5min J band was interrupted because we were reaching low elevations
Marconi QSO0848 Ks 30min we stopped because of twilight - 2.5mag extinction

Details

23:04 Open up the dome - start NB flats
  • observe, flatBrg pisces.20111128.10723.fits - .10763.fits
  • counts varying between 5000counts and 15000counts - cloud passing??

23:35 observe,flatFeII,/no_tcs pisces.20111128.10764.fits - .10844.fits
  • some frames are saturated although we started with counts <19000 - still clouds?

00:15 observe,flatK,/no_tcs pisces.20111128.10845.fits - .10906..fits
  • pisces.20111128.10845.fits - 2sec
  • pisces.20111128.10846.fits - 7sec
  • the rest 8sec
  • there are saturated images 863-881

00:30 observe,flatJ,/no_tcs pisces.20111128.10907.fits - .10918.fits

00:42 starting NB flats for low counts
  • Brg pisces.20111128.10919.fits - .10938.fits
  • FeII pisces.20111128.10939.fits - .10958.fits
  • K pisces.20111128.10959.fits - .11013.fits
  • pisces.20111128.10993.fits - stray light

01:20 we finished with the flats - switched off ventilation/treehouse
  • pointing correction - collimation
  • at this point we also do 'pisces_set_archive' -> next image will be pisces.20111129.0001.fits

01:33 observe, IC10_preset (PA=-180, GS=8)
  • the active optics are collimated - but AO system started with 4micron rms
  • during the first preset CCD39 was blocked - we manually opened the loop and resend the preset
  • the second preset also failed - the loop seemed to be closed but the status of CCD39 was 'not live' - at some point it probably went in time out and the preset was cancelled - the slopes do not arrive from the BCU at the secondary
  • third preset failed again but in another point - the CCD39 is doing what it is supposed to do but nothing arrives at the secondary.
  • from AOS we switch OFF the WFS and start again : Wstop - Wstart
  • from AOS - turn ON
  • during this preset - the loop was closed - slopes started arriving at the secondary but there has been an 'acquire-ref FAIL' - preset cancelled
  • preset again - loop closed with 10 modes and then everything diverges and the mirror resets.

02:30 We restart everything! - AOS reboot

02:39 Call Alfio
  • what happened was: all the system had to be restarted and one subsystem (the arbitrator) had to be restared once more (software bug??)

02:48 problem seems to be solved by switcing OFF-ON the arbitrator - send preset again
  • irc_runao error - but presets seems to have failed before arriving at runao
  • send preset again
  • this time the preset was successful - the loop is closed and stable

03:08 after setting the 0,0 position we start dithering for IC10 in Ks 0008-0017
  • first dither image: pisces.20111129.0008.fits - exptime: 30sec
  • dither on target completed without error

03:17 preset cancelled when it was offsetting 145arcsec (sky image for this target)
  • the offset did not complete
  • offset and obtain sky-image manually - in order to do this we need to resend the preset - close loop - give offset for the sky position (instead of centering the target) and repeat dither
  • the preset was cancelled because the mirror was not yet set - repeat preset
  • the followed procedure (working script) is:
WORKING SCRIPT
# IC10_dither
# 1800sec Ks
header,PARTNER=INAF,OBSERVER='Bono',OBJECT=IC10
set_position
camera, EXP=30.0, FRAMES=2, COADDS=1, FILTER=5
dither, COORD=detxy, TYPE=box, SIZE=2, DIST=1.0, ITER=1,IMAGE=SCIENCE
set_position
camera, EXP=30.0, FRAMES=2, COADDS=1, FILTER=5
dither, COORD=detxy, TYPE=box, SIZE=2, DIST=5.0, ITER=1,IMAGE=SCIENCE
set_position
camera, EXP=30.0, FRAMES=2, COADDS=1, FILTER=5
dither, COORD=detxy, TYPE=box, SIZE=2, DIST=5.0, ITER=1,IMAGE=SCIENCE
FAILING SCRIPT
# IC10_dither
# 1800sec Ks
header,PARTNER=INAF,OBSERVER='Bono',OBJECT=IC10
set_position
camera, EXP=30.0, FRAMES=2, COADDS=1, FILTER=5
dither, COORD=detxy, TYPE=box, SIZE=2, DIST=1.0, ITER=1,IMAGE=SCIENCE
offset,X=145,Y=0,COO=DETXY145
camera, EXP=30.0, FRAMES=2, COADDS=1, FILTER=5
dither, COORD=detxy, TYPE=box, SIZE=2, DIST=5.0, ITER=1,IMAGE=SCIENCE
offset,X=0,Y=0,COO=DETXY
camera, EXP=30.0, FRAMES=2, COADDS=1, FILTER=5
dither, COORD=detxy, TYPE=box, SIZE=2, DIST=5.0, ITER=1,IMAGE=SCIENCE

03:55 preset cancelled - shell ripped for unknown reason
  • repeat preset - cancelled again because the shell was not flat yet...

03:57 the AO software is not responding
  • Adsec-Arbitrator is not responding - stop and restart
  • set the shell - success

04:03 send preset again - IC10_preset.obs
  • guider seeing 1.0'' - 1.2''

04:26 the preset was cancelled - shell ripped
  • it was skipping images and it failed - could it be that seeing was getting worse??
  • dimm seeing 1.6''
  • the AO ref star mag seemed to change - passing cirrus??

04:29 send preset again IC10
  • preset cancelled - it started skipping again

04:39 preset again - the flux is almost 1/3 than previously - probably the clouds are getting thicker

04:44 preset cancelled - too high forces sent at the secondary

04:48 moving to a bright target : T-Tau - send preset
  • 3 presets failed for various reasosn (time out while trying to change bin, the system thought it was in closed loop but it was not...)

05:10 CCD39 went offline (not live)

05:23 send preset - T-Tau
  • dimm seeing 1.1'' - guider seeing 0.5''
  • CCD39 shows high noise
  • CCD39 Sensirion - 14.6C

05:31 after the preset we can see T-Tau - the loop is closed
  • we try to go on with the dither pattern

05:35 Maximum skip reached - AOS error
  • preset failed - try again

06:05 preset was successful - start dithering on T-Tau

07:06 clouds passing by - the AO estimated magnitude is 1.5-2.0mag fainter
  • continue with dithering for as long as loop remains closed
  • dimm seeing 0.8''

07:40 loop opened - send the preset again and go on with the dither

08:25 Repeat T-Tau preset with a small rotation to see if the features we see in the 2.14filter are sources or artifacts
  • iif_offset does not ALWAYS work when you try to insert a rotation because there is not enough time for the rotation to be completed - the loop opens and the preset fails
  • pauses AO - rotates - tries to close the loop again but the mirror does not have enough time to close the loop - it times out and fails
  • it worked when we asked a smaller rotation

08:35 observe,T_Tau with rotation (3-4 angles)

08:42 Slew to NGC1560
  • dimm seeing 0.9''-1.1''
  • pisces.20111129.0498.fits - exptime 2min - the source is in the central position of the dither pattern

09:02 NGC1560 - dither pattern in Ks
  • start series with image pisces.20111129.0500.fits

09:39 the shell ripped - preset cancelled

09:43 repeat NGC1560 preset

09:46 preset cancelled - shell ripped

09:59 loop is closed - send the dither pattern

10:20 shell ripped - we found a HUGE defocus on the guider with the active optics
  • we need to reset the active optics...

10:35 repeat NGC1560 preset - failed because it started AO when active optics rms was ~3000 and it did not manage to close the loop

10:48 repeat NGC1560 preset - shell ripped

10:56 repeat NGC1560 preset - it took 7 min to close the loop

11:03 NGC1560 - dither pattern in J
  • start series with image pisces.20111129.0529.fits

11:14 preset cancelled

11:21 Slew to QSO0848 - no guide star found
  • try again -maybe there is again that big defocus
  • failed - correct pointing

11:30 repeat preset QSO0848
  • after the pointing correction we did find the GS
  • preset failed during runao

11:35 repeat preset QSO0848 - acquire failed - the AO ref star is too faint

11:40 preset QSO0848
  • with the probe on the AO ref star to see if the magnitude is correct
  • we can see the star - we cancel preset

11:43 preset QSO0848
  • first we give an active preset to make it collimate since the star is still too faint for the AO because of abberation
  • then we give the ACE preset QSO0848 dither in Ks
  • example of working script when the initial abberations are too big
header,PARTNER=INAF,OBSERVER='Marconi',OBJECT=QSO0848
preset, OBJECT=QSO0848, MODE=active, GS=1, RS=1, PA=-90
auto_collimate
preset, OBJECT=QSO0848, MODE=ace, GS=6, RS=1, PA=-90
auto_collimate
runao
camera, EXP=1.0, FRAMES=1, COADDS=1, FILTER=5
getimage, IMAGE=SCIENCEheader,PARTNER=INAF,OBSERVER='Marconi',OBJECT=QSO0848
preset, OBJECT=QSO0848, MODE=active, GS=1, RS=1, PA=-90
auto_collimate
preset, OBJECT=QSO0848, MODE=ace, GS=6, RS=1, PA=-90
auto_collimate
runao
camera, EXP=1.0, FRAMES=1, COADDS=1, FILTER=5
getimage, IMAGE=SCIENCE
  • we are getting clouds again

12:05 we are taking images but there are no sources - probably overcast with clouds
  • subtracting two offset images we can see the source
  • we resume with dither pattern

12:15 QSO0848 dither in Ks

12:18 preset cancelled - unknown reason
  • setting the shell - repeat preset

12:32 QSO0848 dither in Ks
  • we get a 3mag extinction - clouds are getting worse...

13:04 preset cancelled - sky too bright and we lost the guide star
  • clouds - no dawn flats

13:08 END OF NIGHT

  • Seeing reported by GCS during AO observations:
    20111129fwhm.jpg

-- KonstantinaBoutsia - 28 Nov 2011
Topic revision: r8 - 25 Jul 2014, DouglasFisher
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback